# Interplay of superconductivity and charge-density-wave order in kagome materials

### v1

@@ III.2 Two-dimensional irrep
$\mathcal{F}^{(2,1,1)}[{\bm{\eta}}_{E_{2}};{\bm{\rho}},\tilde{{\bm{\rho}}}]=\beta_{3}[\frac{\sqrt{3}}{2}(|\eta_{E_{2},1}|^{2}-|\eta_{E_{2},2}|^{2})(\rho_{1}\tilde{\rho}_{1}-\rho_{3}\tilde{\rho}_{3})\\ +(\eta_{E_{2},1}^{\phantom{*}}\eta_{E_{2},2}^{*}+\eta_{E_{2},1}^{*}\eta_{E_{2},2}^{\phantom{*}})(\rho_{2}\tilde{\rho}_{2}-\frac{1}{2}(\rho_{1}\tilde{\rho}_{1}+\rho_{3}\tilde{\rho}_{3}))].$ (22)
@@
Figure 2: The effect of CDW coupling on a two-component order parameter $E_{2}$ with $b_{2}>0$ in Eq. (20). Solid (dashed) lines indicate the component $|\eta_{E_{2},1}|$ ($|\eta_{E_{2},2}|$). For a rotationally symmetric real CDW (blue) with order parameter ${\bm{\rho}}=\rho(1,1,1)$, a $d+id$ order sets in at $T_{\rm c}<T_{\rm c,0}$. If the CDW breaks rotational symmetry (purple), ${\bm{\rho}}=(\rho,\rho+\delta,\rho)^{T}$, the transition is split and a time-reversal symmetric (TRS), anisotropic $d$-wave order is initially preferred. Only at a lower temperature, the second component appears with a phase shift of $\pi/2$. Finally, if the $3{\bm{Q}}$ CDW breaks three-fold rotational symmetry and is imaginary, the CDW again induces a TRSB solution at the transition. $\beta_{2}=0.5$, and $\kappa m_{2}=-0.5$
@@
$\mathcal{F}^{(2,1,1)}[{\bm{\eta}}_{E_{2}};{\bm{\rho}},{\bm{\rho}}^{\prime}]=i\kappa(\eta_{E_{2},1}^{\phantom{*}}\eta_{E_{2},2}^{*}-\eta_{E_{2},1}^{*}\eta_{E_{2},2}^{\phantom{*}})M_{z}.$ (23)
@@
$\mathcal{F}^{(1,1,1)}[{\bm{\eta}}_{E_{2}},{\bm{\eta}};{\bm{\rho}}]=\gamma_{1}\left\{\left[\rho_{2}\eta_{2}-\frac{1}{2}(\rho_{1}\eta_{1}+\rho_{3}\eta_{3})\right]\eta_{E_{2},1}^{*}\right.\\ +\left.\frac{\sqrt{3}}{2}\left(\rho_{1}\eta_{1}-\rho_{3}\eta_{3}\right)\eta_{E_{2},2}^{*}+c.c.\right\}$ (24)
@@
$\mathcal{F}^{(1,1,1)}[{\bm{\eta}}_{E_{2}},{\bm{\eta}};{\bm{\rho}}]=\gamma_{2}\left\{\frac{\sqrt{3}}{2}\left(\rho_{3}\eta_{3}-\rho_{1}\eta_{1}\right)\eta_{E_{2},1}^{*}\right.\\ +\left.\left[\rho_{2}\eta_{2}-\frac{1}{2}(\rho_{1}\eta_{1}+\rho_{3}\eta_{3})\right]\eta_{E_{2},2}^{*}+c.c.\right\}$ (25)
@@ IV.1 PDW in the presence of a CDW
where $a_{\rm PDW}(T)=a_{\rm{PDW},0}(T-T_{\rm c,0})$ changes sign at the critical temperature $T_{\rm c,0}$ and $a_{\rm{PDW},0}>0$. Figure 3 shows the phase diagram for the pure pair density wave [58]. Similar to the case of an $E_{2}$ irrep, here $b_{3}>b_{2}$ leads to a time-reversal-symmetry-breaking state. Note that there are two TRSB pair-density-wave states, namely a chiral $3{\bm{Q}}$ state with ${\bm{\eta}}=|{\bm{\eta}}|(1,e^{i\omega},e^{-i\omega})^{T}$ and $\omega=2\pi/3$ or $\omega=\pi/3$ and a $2{\bm{Q}}$ state with ${\bm{\eta}}=|{\bm{\eta}}|(1,\pm i,0)^{T}$. The latter solution becomes stable for large enough $b_{3}/b_{2}$. In addition, a $3{\bm{Q}}$ real PDW appears for $-b_{3}>b_{2}$ and a $1{\bm{Q}}$ state is realized when the $b_{2}>0$ term dominates, with both orders time-reversal symmetric.
@@
$\mathcal{F}^{(2,1)}[{\bm{\eta}};{\bm{\rho}}]=\frac{\gamma_{\rm PDW}}{2}[\rho_{1}(\eta_{2}^{\phantom{*}}\eta_{3}^{*}+c.c.)+{\rm cyclic}].$ (29)
@@
Figure 4: Crossover from the frustrated regime with $\rho_{i}=\rho$ and $\gamma_{\rm PDW}+\beta_{2}\rho>0$, where the $3{\bm{Q}}$ chiral PDW minimizes the free energy, to a nematic charge density wave, where a PDW with two phases equal to each other and with a $\pi$ difference to the last one is favored. Image obtained with $b_{1}=-b_{2}=1$, $b_{3}=\beta_{1}=0$, $\gamma_{\rm PDW}=0.5$, $\beta_{2}=0.2$, $\beta_{3}=0.1$, and ${\bm{\rho}}\propto(\rho,\rho+\delta,\rho+\delta)^{T}$.
Figure 4 shows the phase diagram for the PDW when the isotropic CDW leads to frustration as a function of CDW anisotropy $\delta$. To not favor a chiral PDW over a real $3{\bm{Q}}$ PDW, we have used $b_{3}=0$, but set $b_{2}<0$ in order to promote a $3{\bm{Q}}$ phase, see Fig. 3. The situation resembles the case of a chiral two-component superconductor, where the anisotropy lifts the degeneracy at $T_{\rm c}$ and thus, the TRSB solution only appears at a temperature below the superconducting transition, see Fig. 2. Note, again, that the chiral solution is not favored by the PDW free energy, but appears due to the frustration of the coupling to the CDW. Intriguingly, a chiral solution even persists for small anisotropy when $b_{3}<0$, in other words when the real $3{\bm{Q}}$ PDW dominates in the pristine phase diagram, Fig. 3. In this special case, the chiral solution is only an intermediate phase that disappears again for small temperatures, where the quartic terms start to dominate.
@@
$\mathcal{F}^{(3)}[{\bm{\eta}};{\bm{\rho}}^{\prime}]=\frac{\gamma_{\rm PDW}^{\prime}}{2}[i\rho_{1}^{\prime}(\eta_{2}^{\phantom{*}}\eta_{3}^{*}-c.c.)+{\rm cyclic}]$ (31)
@@
$\mathcal{F}^{(2,1,1)}[{\bm{\eta}};{\bm{\rho}},{\bm{\rho}}^{\prime}]$ $=i\frac{\beta_{4}}{2}[(\rho_{2}\rho_{3}^{\prime}+\rho_{2}^{\prime}\rho_{3})(\eta_{2}^{\phantom{*}}\eta_{3}^{*}-c.c.)$ (32)
@@ IV.2 PDW as the only instability
For completeness, we finish our considerations with the situation, where the leading instability is only towards a PDW. This situation has been discussed in Ref. 90 and for AVS in Ref. 58.
@@ Appendix A Extended point group $C_{6v}^{\prime\prime\prime}$
In this appendix, we provide the character table for the extended point group $C_{6v}^{\prime\prime\prime}$, Tab. 2, as well as the decomposition of the symmetrized products for the three-dimensional irreps, Tab. 3. The full product table for the irreps of $C_{6v}^{\prime\prime\prime}$ can be found, e.g., in Ref. 67.
@@ Appendix B Visualization of pair-density-wave orders
$F_{1}$ $F_{1,1}$ $F_{1,2}$ $F_{1,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}$ $F_{3,1}$ $F_{3,2}$ $F_{3,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}$ $F_{4,1}$ $F_{4,2}$ $F_{4,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 4: Real space illustrations of on-site pairing states on the kagome lattice which break translational symmetry. The color (red/blue) denotes the sign ($+1$/$-1$) of the on-site singlet pairing interaction. The first column provides the three-dimensional irrep, denoted according to the $C_{6v}^{\prime\prime\prime}$ point group. The second column shows each component of the irrep and the last column shows a symmetric superposition of the three components. The black parallelogram illustrates the $2\times 2$ enlarged unit cell.
@@
$F_{1}^{(1)}$ $F_{1,1}^{(1)}$ $F_{1,2}^{(1)}$ $F_{1,3}^{(1)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{1}^{(2)}$ $F_{1,1}^{(2)}$ $F_{1,2}^{(2)}$ $F_{1,3}^{(2)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{2}$ $F_{2,1}$ $F_{2,2}$ $F_{2,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}^{(1)}$ $F_{3,1}^{(1)}$ $F_{3,2}^{(1)}$ $F_{3,3}^{(1)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}^{(2)}$ $F_{3,1}^{(2)}$ $F_{3,2}^{(2)}$ $F_{3,3}^{(2)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}$ $F_{4,1}$ $F_{4,2}$ $F_{4,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 5: Real space illustrations of nearest-neighbor singlet pairing states on the kagome lattice which break translational symmetry. The color (red/blue) denotes the sign ($+1$/$-1$) of the singlet pairing interaction. The first column provides the three-dimensional irrep, denoted according to the $C_{6v}^{\prime\prime\prime}$ point group. The second column shows each component of the irrep and the last column shows a symmetric superposition of the three components. The black parallelogram illustrates the $2\times 2$ enlarged unit cell.
@@
$F_{1}$ $F_{1,1}$ $F_{1,2}$ $F_{1,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{2}^{(1)}$ $F_{2,1}^{(1)}$ $F_{2,2}^{(1)}$ $F_{2,3}^{(1)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{2}^{(2)}$ $F_{2,1}^{(2)}$ $F_{2,2}^{(2)}$ $F_{2,3}^{(2)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}$ $F_{3,1}$ $F_{3,2}$ $F_{3,3}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}^{(1)}$ $F_{4,1}^{(1)}$ $F_{4,2}^{(1)}$ $F_{4,3}^{(1)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}^{(2)}$ $F_{4,1}^{(2)}$ $F_{4,2}^{(2)}$ $F_{4,3}^{(2)}$ [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Real space illustrations of nearest-neighbor triplet pairing states on the kagome lattice which break translational symmetry. The first column provides the three-dimensional irrep, denoted according to the $C_{6v}^{\prime\prime\prime}$ point group. The second column shows each component of the irrep and the last column shows a symmetric superposition of the three components. The black parallelogram illustrates the $2\times 2$ enlarged unit cell.

### v2

@@ III.2 Two-dimensional irrep
$\mathcal{F}^{(2;1,1)}[{\bm{\eta}}_{E_{2}};{\bm{\rho}},\tilde{{\bm{\rho}}}]=\beta_{3}[\frac{\sqrt{3}}{2}(|\eta_{E_{2},1}|^{2}-|\eta_{E_{2},2}|^{2})(\rho_{1}\tilde{\rho}_{1}-\rho_{3}\tilde{\rho}_{3})\\ +(\eta_{E_{2},1}^{\phantom{*}}\eta_{E_{2},2}^{*}+\eta_{E_{2},1}^{*}\eta_{E_{2},2}^{\phantom{*}})(\rho_{2}\tilde{\rho}_{2}-\frac{1}{2}(\rho_{1}\tilde{\rho}_{1}+\rho_{3}\tilde{\rho}_{3}))].$ (22)
@@
Figure 5: The effect of CDW coupling on a two-component order parameter $E_{2}$ with $b_{2}>0$ in Eq. (20). Solid (dashed) lines indicate the component $|\eta_{E_{2},1}|$ ($|\eta_{E_{2},2}|$). For a rotationally symmetric real CDW (blue) with order parameter ${\bm{\rho}}=\rho(1,1,1)$, a $d+id$ order sets in at $T_{\rm c}<T_{\rm c,0}$. If the CDW breaks rotational symmetry (purple), ${\bm{\rho}}=(\rho,\rho+\delta,\rho)^{T}$, the transition is split and a time-reversal symmetric (TRS), anisotropic $d$-wave order is initially preferred. Only at a lower temperature, the second component appears with a phase shift of $\pi/2$. Finally, if the $3{\bm{Q}}$ CDW breaks three-fold rotational symmetry and is imaginary, the CDW again induces a TRSB solution at the transition. Parameters are chosen as $a_{E_{2},0}=T_{\rm c,0}=b_{1}=1$, $b_{2}=0.8$, $\beta_{1}=0.025$, $\beta_{2}=-0.5$, $\kappa m_{2}=-0.5$, $\delta/\rho=2/7$, and $|\bm{\rho}|=1$.
@@
$\mathcal{F}^{(2;1,1)}[{\bm{\eta}}_{E_{2}};{\bm{\rho}},{\bm{\rho}}^{\prime}]=i\kappa(\eta_{E_{2},1}^{\phantom{*}}\eta_{E_{2},2}^{*}-\eta_{E_{2},1}^{*}\eta_{E_{2},2}^{\phantom{*}})M_{z}.$ (23)
@@
$\mathcal{F}^{(1,1;1)}[{\bm{\eta}}_{E_{2}},{\bm{\eta}};{\bm{\rho}}]=\gamma_{1}\left\{\left[\rho_{2}\eta_{2}-\frac{1}{2}(\rho_{1}\eta_{1}+\rho_{3}\eta_{3})\right]\eta_{E_{2},1}^{*}\right.\\ +\left.\frac{\sqrt{3}}{2}\left(\rho_{1}\eta_{1}-\rho_{3}\eta_{3}\right)\eta_{E_{2},2}^{*}+c.c.\right\}$ (24)
@@
$\mathcal{F}^{(1,1;1)}[{\bm{\eta}}_{E_{2}},{\bm{\eta}};{\bm{\rho}}]=\gamma_{2}\left\{\frac{\sqrt{3}}{2}\left(\rho_{3}\eta_{3}-\rho_{1}\eta_{1}\right)\eta_{E_{2},1}^{*}\right.\\ +\left.\left[\rho_{2}\eta_{2}-\frac{1}{2}(\rho_{1}\eta_{1}+\rho_{3}\eta_{3})\right]\eta_{E_{2},2}^{*}+c.c.\right\}$ (25)
@@ IV.1 PDW in the presence of a CDW
where $a_{\rm PDW}(T)=a_{\rm{PDW},0}(T-T_{\rm c,0})$ changes sign at the critical temperature $T_{\rm c,0}$ and $a_{\rm{PDW},0}>0$ and the $b_{i}$ are temperature-independent phenomenological parameters. Figure 6 shows the phase diagram for the pure pair density wave [61]. Similar to the case of an $E_{2}$ irrep, here $b_{3}>b_{2}$ leads to a time-reversal-symmetry-breaking state. Note that there are two TRSB pair-density-wave states, namely a chiral $3{\bm{Q}}$ state with ${\bm{\eta}}=|{\bm{\eta}}|(1,e^{i\omega},e^{-i\omega})^{T}$ and $\omega=2\pi/3$ or $\omega=\pi/3$ and a $2{\bm{Q}}$ state with ${\bm{\eta}}=|{\bm{\eta}}|(1,\pm i,0)^{T}$. The latter solution becomes stable for large enough $b_{3}/b_{2}$. In addition, a $3{\bm{Q}}$ real PDW appears for $-b_{3}>b_{2}$ and a $1{\bm{Q}}$ state is realized when the $b_{2}>0$ term dominates, with both orders time-reversal symmetric. The real space phase structure of the PDW order parameters is illustrated in Fig. 7.
@@
$\mathcal{F}^{(2;1)}[{\bm{\eta}};{\bm{\rho}}]=\frac{\gamma_{\rm PDW}}{2}[\rho_{1}(\eta_{2}^{\phantom{*}}\eta_{3}^{*}+c.c.)+{\rm cyclic}].$ (29)
@@
Figure 8: Phase diagram showing the transition from the frustrated regime with $\rho_{i}=\rho$ and $\gamma_{\rm PDW}+\beta_{2}\rho>0$, where the $3{\bm{Q}}$ chiral PDW minimizes the free energy, to a nematic charge density wave, where a PDW with two phases equal to each other and with a $\pi$ difference to the last one is favored. Figure obtained with $a_{\rm PDW,0}=T_{\rm c,0}=b_{1}=-b_{2}=1$, $b_{3}=\beta_{1}=0$, $\gamma_{\rm PDW}=0.5$, $\beta_{2}=0.2$, $\beta_{3}=0.1$, ${\bm{\rho}}\propto(\rho,\rho+\delta,\rho+\delta)^{T}$, and $|\bm{\rho}|=1$. The dashed black line shows the TRSB-TRS transition for $b_{3}=-0.1$.
Figure 8 shows the phase diagram for the PDW when the isotropic CDW leads to frustration as a function of CDW anisotropy $\delta$. To not favor a chiral PDW over a real $3{\bm{Q}}$ PDW, we have used $b_{3}=0$, but set $b_{2}<0$ in order to promote a $3{\bm{Q}}$ phase, see Fig. 6. While the critical temperature of the PDW generically changes when changing the CDW order parameter, we have further kept this change minimal by fixing the absolute value $|\bm{\rho}|$. The situation described in Fig. 8 resembles the case of a chiral two-component superconductor, where the anisotropy lifts the degeneracy at $T_{\rm c}$ and thus, the TRSB solution only appears at a temperature below the superconducting transition, see Fig. 5. Note, again, that the chiral solution is not favored by the PDW free energy, but appears due to the frustration of the coupling to the CDW. Intriguingly, a chiral solution even persists for small anisotropy when $b_{3}<0$, in other words when the real $3{\bm{Q}}$ PDW dominates in the pristine phase diagram, Fig. 6. In this special case, the chiral solution is only an intermediate phase that disappears again for small temperatures, where the quartic terms start to dominate.
@@
$\mathcal{F}^{(2;1)}[{\bm{\eta}};{\bm{\rho}}^{\prime}]=\frac{\gamma_{\rm PDW}^{\prime}}{2}[i\rho_{1}^{\prime}(\eta_{2}^{\phantom{*}}\eta_{3}^{*}-c.c.)+{\rm cyclic}]$ (31)
@@
$\mathcal{F}^{(2;1,1)}[{\bm{\eta}};{\bm{\rho}},{\bm{\rho}}^{\prime}]$ $=i\frac{\beta_{4}}{2}[(\rho_{2}\rho_{3}^{\prime}+\rho_{2}^{\prime}\rho_{3})(\eta_{2}^{\phantom{*}}\eta_{3}^{*}-c.c.)$ (32)
@@ IV.2 PDW as the only instability
For completeness, we finish our considerations with the situation, where the leading instability is only towards a PDW. This situation has been discussed in Ref. 94 and for AVS in Ref. 61.
@@ Appendix A Extended point group $C_{6v}^{\prime\prime\prime}$
In this appendix, we provide the character table for the extended point group $C_{6v}^{\prime\prime\prime}$, Tab. 2, as well as the decomposition of the symmetrized products for the three-dimensional irreps, Tab. 3. The full product table for the irreps of $C_{6v}^{\prime\prime\prime}$ can be found, e.g., in Ref. 70.
@@ Appendix B Visualization of pair-density-wave orders
$F_{1}$ $F_{1,1}$ $F_{1,2}$ $F_{1,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}$ $F_{3,1}$ $F_{3,2}$ $F_{3,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}$ $F_{4,1}$ $F_{4,2}$ $F_{4,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 4: Real space illustrations of on-site pairing states on the kagome lattice which break translational symmetry. The color (red/blue) denotes the sign ($+1$/$-1$) of the on-site singlet pairing interaction. The first column provides the three-dimensional irrep, denoted according to the $C_{6v}^{\prime\prime\prime}$ point group. The second column shows each component of the irrep, and the last column shows a symmetric superposition of the three components.
@@
$F_{1}^{(1)}$ $F_{1,1}^{(1)}$ $F_{1,2}^{(1)}$ $F_{1,3}^{(1)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{1}^{(2)}$ $F_{1,1}^{(2)}$ $F_{1,2}^{(2)}$ $F_{1,3}^{(2)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{2}$ $F_{2,1}$ $F_{2,2}$ $F_{2,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}^{(1)}$ $F_{3,1}^{(1)}$ $F_{3,2}^{(1)}$ $F_{3,3}^{(1)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}^{(2)}$ $F_{3,1}^{(2)}$ $F_{3,2}^{(2)}$ $F_{3,3}^{(2)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}$ $F_{4,1}$ $F_{4,2}$ $F_{4,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 5: Real space illustrations of nearest-neighbor singlet pairing states on the kagome lattice which break translational symmetry. The color (red/blue) denotes the sign ($+1$/$-1$) of the singlet pairing interaction. The first column provides the three-dimensional irrep, denoted according to the $C_{6v}^{\prime\prime\prime}$ point group. The second column shows each component of the irrep, and the last column shows a symmetric superposition of the three components.
@@
$F_{1}$ $F_{1,1}$ $F_{1,2}$ $F_{1,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{2}^{(1)}$ $F_{2,1}^{(1)}$ $F_{2,2}^{(1)}$ $F_{2,3}^{(1)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{2}^{(2)}$ $F_{2,1}^{(2)}$ $F_{2,2}^{(2)}$ $F_{2,3}^{(2)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{3}$ $F_{3,1}$ $F_{3,2}$ $F_{3,3}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}^{(1)}$ $F_{4,1}^{(1)}$ $F_{4,2}^{(1)}$ $F_{4,3}^{(1)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$F_{4}^{(2)}$ $F_{4,1}^{(2)}$ $F_{4,2}^{(2)}$ $F_{4,3}^{(2)}$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Real space illustrations of nearest-neighbor triplet pairing states on the kagome lattice which break translational symmetry. The first column provides the three-dimensional irrep, denoted according to the $C_{6v}^{\prime\prime\prime}$ point group. The second column shows each component of the irrep, and the last column shows a symmetric superposition of the three components.